OpenStudy (anonymous):

solve the system by elimination -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5 I don't understand this at all if someone could come up with a full explanation ill greatly appreciate it.

OpenStudy (jdoe0001):

do you know how to solve by elimination with 2 variables?

OpenStudy (anonymous):

Im sorry i barely understand this

OpenStudy (jdoe0001):

well... do you even know what solving by elimination mean?

OpenStudy (anonymous):

yes i do

OpenStudy (jdoe0001):

have you done any with only 2 variables?

OpenStudy (anonymous):

yes i dont know how to do it with the three variables

OpenStudy (jdoe0001):

the way it works is, off the 3 equations, you grab 2 at a time a time, and eliminate 1 variable then you go and grab 2 more and eliminate same 1 variable -2x+2y+3z=0 <---- let's use this one -2x-y+z=-3 <---- and this one 2x+3y+3z=5 \(\begin{array}{llll} -2x+2y+3z=0&\implies &-2x+2y+3z=0\\ -2x-y+z=-3& \times -3\implies &6x+3y-3z=9\\ \hline\\ &&4x+5y+0=9\\ \quad \\ \color{blue}{4x+5y=9} \end{array}\)

OpenStudy (jdoe0001):

now let's use grab another set of 2.... let's grab this time -2x+2y+3z=0 <--- this one again -2x-y+z=-3 2x+3y+3z=5 <--- and this one this time \(\begin{array}{llll} -2x+2y+3z=0&\implies &-2x+2y+3z=0\\ 2x+3y+3z=5& \times -1\implies &-2x-3y-3z=-5\\ \hline\\ &&-4x-y+0=-5\\ \quad \\ \color{blue}{-4x-y=-5} \end{array}\)

OpenStudy (jdoe0001):

and now you just grab those 2 RESULTANT equations, with only 2 variables and solve for either, then get the other, and then plug both in to get "z" in any of the 3 equations above

OpenStudy (anonymous):

thank you i sort of understand it now

OpenStudy (anonymous):

i got x=1.5 y=-1 z=1.6 is that right?